OpenStudy (atomicreaper):

The sum of the squares of two positive integers is 185. If one integer is 3 less than the other, Find the larger integer.

OpenStudy (welshfella):

if we let highest number be x then other number is x - 3 so we have x^2 + (x -3)^2 = 185 solve this for x.

OpenStudy (atomicreaper):

@welshfella, would it be 11?

OpenStudy (welshfella):

yes

OpenStudy (atomicreaper):

Ok thanks!!!

OpenStudy (atomicreaper):

Do you mind helping me set up another problem?

OpenStudy (welshfella):

the eqtion simplifies to x^2 - 3x - 88 = 0 (x - 11)(x - 8) = 0

OpenStudy (welshfella):

I have 10 minites before i go

OpenStudy (atomicreaper):

I just need help setting it up.

OpenStudy (welshfella):

ok

OpenStudy (welshfella):

that last should have been (x - 11)(x + 8) = 0

OpenStudy (atomicreaper):

I tagged you in the new one so i could give another medal